OpenStudy (anonymous):

Show that: [(1/(x+h)) - (1/x)]/h = -1/(x(x+h))

OpenStudy (nincompoop):

show it

geerky42 (geerky42):

Let me clear things up: \[\dfrac{\left(\dfrac{1}{x+h} - \dfrac{1}{x}\right)}{h} = \dfrac{-1}{x(x+h)}\]

geerky42 (geerky42):

Hmm, I would get denominator same if I were you

OpenStudy (triciaal):

to show do each side separate and simplify you will see at the very end they are equal and therefore you have shown that they are equal. for this problem left hand side